OpenStudy (anonymous):

find the LCM of 10 and 12

OpenStudy (anonymous):

For example, for LCM (12,30) we find: Prime factorization of 12 = 2 * 2 * 3 = 22 * 31 * 5. Prime factorization of 30 = 2 * 3 * 5 = 21 * 31 * 5. Using the set of prime numbers from each set with the highest exponent value we take 22 * 31 * 51 = 60.